A Sociology instructor gives students points for each​ discussion-board post and points for each reply to a post. Ana wrote 5 posts and 8 replies and received 41 points. Jae wrote 3 posts and 4 replies and received 23 points. Determine how many points a discussion post is worth and how many points a reply is worth.
Answer by Boreal(15235)

5p+8r=41
3p+4r=23
eliminate r by multiplying the second equation by -2 and adding the two.
-6p-8r=-46
-p=-5
p=5 points per post
r=2 points per reply